# Training-Free Label Space Alignment for Universal Domain Adaptation

### v1

@@ 5 Results
Table 3: H-score (%) and H3-score (%) comparisons between existing methods and the proposed method using the CLIP ViT-B/16 backbone across four UniDA settings (open-partial, open, closed, partial). An asterisk (*) represents results for some methods referenced from the previous paper Lai et al. (2023). A dash (-) represents that any metrics not reported in the official papers.
@@ 5.1 Quantitative Comparison
Comparison with baseline methods
We compared TLSA against existing UniDA approaches. Among these, CLIP-Distillation achieves state-of-the-art (SOTA) performance. Except for the Office31 open-set and VisDA partial settings, we attain SOTA across all scenarios. Although the impact of self-training is more significant on certain datasets, it generally enhances the H3-score in most cases. The results are summarized in Table 3. Since calculating the H-score and H3-score for partial and closed-set scenarios is infeasible, we employed the average accuracy score as a substitute for the H-score and H3-score, following Deng and Jia (2023). These findings highlight the versatility and effectiveness of TLSA and TLSA+ST (Self-Training) across diverse UniDA settings, particularly in open-partial and open scenarios where managing unseen classes is critical.
@@
Tables 4 and 5 compare our results with existing methods in the open partial domain adaptation scenario using both ResNet50 (ImageNet-pretrained) and ViT-B/16 (CLIP-pretrained) backbones. Our approach achieved state-of-the-art performance across nearly all domain-to-domain adaptations. However, a slight reduction in H-score was observed due to self-training in DomainNet. While closed-set accuracy improved, OOD detection performance tended to decrease, and vice versa. This indicates a trade-off: self-training shifts the focus of the model, often leading to a decline in H-score performance.
Robustness in various UniDA
@@
Figure 6: Robustness across different class split settings in Office-Home, RealWorld to Clipart, and DomainNet Real to Painting.
Table 6: List of ground truth target private labels and final candidates extracted from TLSA for the Office31, dslr, open-partial setting. The same color indicates the same semantic group.
True target private Predicted target private
@@ 5.2 Qualitative Comparison
H-score: 41.7%
(a) CLIP zero-shot
H-score: 66.02%
(b) TLSA+ST ($|C_{s}|{+}1$ classifier)
H-score: 90.2%
(c) TLSA+ST (Universal classifier)
[…]
Feature visualization
We present t-SNE (Long et al., 2016) plots in Fig. 7 to compare feature discrimination between the CLIP zero-shot and $|C_{s}|+1$classifiers using our approach. As other methods employ a fixed backbone identical to that in (Deng and Jia, 2023), we omitted their representations from the visualization. Furthermore, to emphasize our universal classifier’s discriminative ability, we utilized a $|C_{s}|+1$classifier with LP-ST (linear probing via self-distillation and self-training using top-$k$ confident sample selection), similar to LP-FT (Kumar et al., 2022). The $|C_{s}|+1$-way classifier underperforms compared to TLSA due to its constrained capacity to represent unknown classes.
Predicted target-private labels analysis
Table 6 presents examples of the target-private set extracted by the proposed method. We observed that these predicted labels align with the ground-truth target-private categories in the target domain under the DSLR setting of Office31, indicating that TLSA effectively identifies genuine target-private classes. Importantly, the discovered labels do not need to precisely match the ground-truth annotations; they may also capture prominent attributes of target-domain images. This flexibility is beneficial because generative VLMs can function effectively even on objects not encountered during pretraining, making them particularly suited for open vocabulary or unknown classes. This process generates a set of discovered labels, $R=\{r_{i}\}_{i=1}^{N}$, where $N$ represents the batch size of the target image.
Analysis of the number of predicted target private labels
Fig. 8 displays the number of predicted target-private labels in our study. Across various datasets and settings, our method accurately predicts target-private labels compared to their ground-truth counterparts. This illustrates the flexibility of our design in adapting to different split configurations and datasets.
Figure 8: The number of predicted target private labels through change of split settings and datasets.
@@ 5.3 Ablation Study
Table 7: Ablation study with different components of our alignment pipeline approach. The semantic alignment strategy refers to the comparison of alignment methods used in step 2. Top-1 refers to top-1 prediction-based alignment, Margin refers to margin-based alignment, and Align refers to our method.
@@
Table 8: Ablation study comparing H-scores across different prediction set decision thresholds.
@@
Ablation Study and Comparison of the Alignment Pipeline
We implement an ablation study of label space alignment. As can be seen in Table 7, Synonym label alignment serves as a fundamental filtering method, contributing to performance to a certain extent.
Frequency-based noisy candidate filtering is a vital component of our pipeline; without it, highly noisy candidate labels could be incorrectly designated as target-private labels. A comparison of these methods is shown in Table 7.
Ablation Study of Two Criteria for Label Space Alignment
We conducted an ablation study on the two thresholds, $\tau_{\text{avg}}$and $\tau_{\text{gap}}$, used in adaptive filtering. As outlined in Table 8, $\tau_{\text{avg}}$significantly impacts performance more than $\tau_{\text{gap}}$. Furthermore, $\tau_{\text{avg}}$and $\tau_{\text{gap}}$operate as complementary thresholds. Our design ensures robustness against split-setting variations and diverse datasets.
@@
Ablation Study of Self-Training Components
We conducted experiments to enhance the reliability of our self-training design. The key observation was that incorporating delayed EMA components (Delay in Table 9) slowed the training process of the teacher model and significantly improved the model performance. Moreover, our approach surpassed the top-k% pseudo-label sampling strategy proposed in a previous study (Zara et al., 2023) (Table  9).
5.4 Further Analysis
@@
Sensitivity Analysis for Hyperparameters in TLSA
We analyze the sensitivity of our hyperparameters, $k$, in the top-$k$ prediction within the label-space alignment pipeline and $\epsilon$ in frequency-based noisy candidate filtering. The results are presented in Table 10. The experiments on the DomainNet dataset determined the optimal value based on average performance.
@@
Scalability Analysis for Model Sizes
We conducted experiments using the CLIP-ViT-B/16 backbone. A crucial aspect of validating our method is determining whether performance scales with the backbone size. As illustrated in Table  11, increasing the backbone to CLIP-ViT-L/14@336px leads to a significant performance enhancement.
Learning Statistics of Self-Training
We present self-training learning statistics for Office31 DSLR and Office-Home Clipart under the OPDA setting in Fig. 9. A notable performance improvement occurs after a certain number of iterations.
Figure 9: Learning statistics of our self-training. The x-axis value represents the iteration.
Trade-off for training using the source samples.
Table 12: Comparison of H-scores between self-training combined with source training and ours (self-training only).
@@
We observe that incorporating the source loss term $L_{src}$ introduces a tradeoff. Table 12 demonstrates that integrating source training significantly boosts performance in partial domain adaptation (PDA) and closed-set domain adaptation (CDA) compared to self-training. However, this approach leads to higher computational demands and slightly reduced performance in open-set domain adaptation (OPDA) and open domain adaptation (ODA). Therefore, the method can be selectively employed based on the objective (Table  12).
@@ 6 Limitations
This study has two key limitations. First, it may not be suitable for fine-grained datasets. Because it depends on the zero-shot capabilities of CLIP and BLIP, classes that are fine-grained and underrepresented during pretraining are prone to inadequate recognition, as highlighted by Zhao et al. (2024). Furthermore, the approach does not individually address source-private class detection, which may lead to less than optimal performance within the universal set. We anticipate that our research will advance future investigations in this domain.
@@ 7 Conclusion
In this paper, we introduce a TLSA, a framework leveraging CLIP to address the UniDA problem through label space alignment. This framework utilizes a generative vision-language model to discover labels for all target domain samples and then filter out those containing source-related information. To achieve this, on a mini-batch basis, we remove trivial synonyms of source labels by synonym label alignment and remove semantically ambiguous similar labels by semantic label alignment. Subsequently, on a dataset-wide scale, frequency-based noisy candidate filtering identifies and removes unreliable labels based on occurrence patterns. The method is highly efficient both temporally and memory-wise, as it completes the entire process in a single epoch without the need for further tuning. After completing all alignment processes, we extend the existing source classes with the newly predicted target private labels to create a new universal classifier. This classifier demonstrates significantly better performance compared to existing methods. Additionally, by self-training with balanced top-k confident sample selection, we surpass the state-of-the-art UniDA baselines.
Appendix
8 Pseudo code
@@ 9 Examples of label space alignment
10 Details of semantic label alignment
@@
11 Analysis of labels to be added to the frequency bank
In Section 3.4, we noted that storing BLIP’s discovered label to the frequency bank is more beneficial for performance compared to storing CLIP’s top-1 prediction to the frequency bank. As shown in Table 13, storing CLIP’s top-1 predictions often leads to an overestimation of incorrect target private labels (e.g., overly specific or entirely incorrect classes), which in turn causes the correct target private labels (e.g., ‘train’, ‘truck’, ‘skateboard’) to be underestimated. This imbalance ultimately results in degraded model performance. In Table 13, correct target private label is underestimated when using CLIP’s top-1 prediction compared to using BLIP’s discovered label.
@@ 11 Analysis of labels to be added to the frequency bank
12 Details of prompts to generative VLM
This section provides details of the question prompts used for label discovery. To ensure diversity while maintaining the same meaning, we employed five different styles of prompts. Table 14 demonstrates the detailed examples of question prompts for generative VLM.
@@ 12 Details of prompts to generative VLM
13 Time complexity of our method
This section explains the time complexity of our method. In Fig. 12, despite utilizing two VLMs, our approach adopts a training-free framework, resulting in minimal overhead compared to existing works. Moreover, it demonstrates faster adaptation than clustering-based methods (Chang et al., 2022).
@@ 13 Time complexity of our method
14 Details of main results
This section details the experimental results (H-score, H3-score) for domain-to-domain. Table 16 through Table 26 offer a comprehensive comparison between baseline methods and the proposed approach using the CLIP ViT-B/16 backbone.

### v2

@@ 5 Results
Table 3: H-score (%) and H3-score (%) comparisons between existing methods and the proposed method using the CLIP ViT-B/16 backbone across four UniDA settings (open-partial, open, closed, partial). An asterisk (*) represents results for some methods referenced from the previous paper (Lai et al., 2023). A dash (-) represents that any metrics not reported in the official papers.
@@ 5.1 Quantitative Comparison
Comparison with Baseline Methods. We compared TLSA against existing UniDA approaches. Among these, CLIP-Distillation achieves state-of-the-art (SOTA) performance. Except for the Office31 open-set and VisDA partial settings, we attain SOTA across all scenarios. Although the impact of self-training is more significant on certain datasets, it generally enhances the H3-score in most cases. The results are summarized in Tab. 3. Since calculating the H-score and H3-score for partial and closed-set scenarios is infeasible, we employed the average accuracy score as a substitute for the H-score and H3-score, following Deng and Jia (2023). These findings highlight the versatility and effectiveness of TLSA and TLSA+ST (Self-Training) across diverse UniDA settings, particularly in open-partial and open scenarios where managing unseen classes is critical.
@@
Figure 6: Robustness across different class split settings in Office-Home (left) RealWorld to Clipart, and DomainNet (right) Real to Painting.
Figure 7: The number of predicted target-private labels across different split settings and datasets.
Tab. 4 and 5 compare our results with existing methods in the open partial domain adaptation scenario using both ResNet50 (ImageNet-pretrained) and ViT-B/16 (CLIP-pretrained) backbones. Our approach achieved state-of-the-art performance across nearly all domain-to-domain adaptations. However, a slight reduction in H-score was observed due to self-training in DomainNet. While closed-set accuracy improved, OOD detection performance tended to decrease, and vice versa. This indicates a trade-off: self-training shifts the focus of the model, often leading to a decline in H-score performance.
@@
H-score: 41.7%
(a) CLIP zero-shot
H-score: 66.02%
(b) TLSA+ST ($|C_{s}|{+}1$ classifier)
H-score: 90.2%
(c) TLSA+ST (Universal classifier)
Figure 8: The t-SNE plots of target domain samples for the zero-shot and TLSA+ST methods in the Office-Home RealWorld $\to$ Clipart OPDA setting. Blue points indicate target-private samples, while red points represent source-class samples.
Table 6: List of ground truth target-private labels and final candidates extracted from TLSA for the Office31, DSLR, open-partial setting. The same color indicates the same semantic group.
True target-private Predicted target-private
@@ 5.2 Qualitative Comparison
Analysis of the Number of Predicted Target-Private Labels. Fig. 7 displays the number of predicted target-private labels in our study. Across various datasets and settings, our method accurately predicts target-private labels compared to their ground-truth counterparts. This illustrates the flexibility of our design in adapting to different split configurations and datasets.
Feature Visualization. We present t-SNE (Long et al., 2016) plots in Fig. 8 to compare feature discrimination between the CLIP zero-shot and $|C_{s}|+1$classifiers using our approach. As other methods employ a fixed backbone identical to that in (Deng and Jia, 2023), we omitted their representations from the visualization. Furthermore, to emphasize our universal classifier’s discriminative ability, we utilized a $|C_{s}|+1$classifier with LP-ST (linear probing via self-distillation and self-training using top-$k$ confident sample selection), similar to LP-FT (Kumar et al., 2022). The $|C_{s}|+1$-way classifier underperforms compared to TLSA due to its constrained capacity to represent unknown classes.
Predicted Target-Private Labels Analysis. Tab. 6 presents examples of the target-private set extracted by the proposed method. We observed that these predicted labels align with the ground-truth target-private categories in the target domain under the DSLR setting of Office31, indicating that TLSA effectively identifies genuine target-private classes. Importantly, the discovered labels do not need to precisely match the ground-truth annotations; they often also capture prominent attributes of target-domain images. This flexibility is beneficial because generative VLMs can function effectively even on objects not encountered during pretraining, making them particularly suited for open vocabulary or unknown classes. This process generates a set of discovered labels, $R=\{r_{i}\}_{i=1}^{N}$, where $N$ represents the batch size of the target image.
@@ 5.3 Ablation Study
Ablation Study and Comparison of the Alignment Pipeline. We implement an ablation study of label space alignment. As can be seen in Tab. 7, Synonym label alignment serves as a fundamental filtering method, contributing to performance to a certain extent.
Frequency-based noisy candidate filtering is a vital component of our pipeline; without it, highly noisy candidate labels could be incorrectly designated as target-private labels. A comparison of these methods is shown in Tab. 7.
Table 7: Ablation study with different components of our alignment pipeline approach. The semantic alignment strategy refers to the comparison of alignment methods used in step 2. Top-1 refers to top-1 prediction-based alignment, Margin refers to margin-based alignment, and TLSA refers to our method.
@@
Table 8: Ablation study comparing H-scores across different prediction set decision thresholds
@@
Ablation Study of Two Criteria for Label Space Alignment. We conducted an ablation study on the two thresholds, $\tau_{\text{avg}}$and $\tau_{\text{gap}}$, used in adaptive filtering. As outlined in Tab. 8, $\tau_{\text{avg}}$significantly impacts performance more than $\tau_{\text{gap}}$. Furthermore, $\tau_{\text{avg}}$and $\tau_{\text{gap}}$operate as complementary thresholds. Our design ensures robustness against split-setting variations and diverse datasets.
Ablation Study of Self-Training Components. We conducted experiments to enhance the reliability of our self-training design. The key observation was that incorporating delayed EMA components (Delay in Tab. 9) slowed the training process of the teacher model and significantly improved the model performance. Moreover, our approach surpassed the top-k% pseudo-label sampling strategy proposed in a previous study (Zara et al., 2023) (Tab.  9).
5.4 Further Analysis
@@
Sensitivity Analysis for Hyperparameters in TLSA. We analyze the sensitivity of our hyperparameters, $k$, in the top-$k$ prediction within the label space alignment pipeline and $\epsilon$ in frequency-based noisy candidate filtering. The results are presented in Tab. 10. The experiments on the DomainNet dataset determined the optimal value based on average performance.
@@
Scalability Analysis for Model Sizes. We conducted experiments using the CLIP-ViT-B/16 backbone. A crucial aspect of validating our method is determining whether performance scales with the backbone size. As illustrated in Tab. 11, increasing the backbone to CLIP-ViT-L/14@336px leads to a significant performance enhancement.
Figure 9: Learning Statistics of Our Self-Training. The x-axis denotes the iteration.
Learning Statistics of Self-Training. We present self-training learning statistics for Office31 DSLR and Office-Home Clipart under the OPDA setting in Fig. 9. A notable performance improvement occurs after a certain number of iterations.
Table 12: Comparison of H-scores between self-training combined with source training and ours (self-training only)
@@
Trade-off for Training Using the Source Samples. We observe that incorporating the source loss term $L_{src}$ introduces a tradeoff. Tab. 12 demonstrates that integrating source training significantly boosts performance in partial domain adaptation (PDA) and closed-set domain adaptation (CDA) compared to self-training. However, this approach leads to higher computational demands and slightly reduced performance in open-set domain adaptation (OPDA) and open domain adaptation (ODA). Therefore, the method can be selectively employed based on the objective (Tab. 12).
@@ 6 Limitations
A key limitation of our approach is that it does not explicitly address source-private class detection, which may result in suboptimal performance within the universal label space. However, we note that explicit source-private class detection often leads to mis-detection in practice, and focusing on target-private class detection can achieve most of the performance gains.
@@ 7 Conclusion
In this paper, we introduce Training-free Label Space Alignment (TLSA), a framework that leverages CLIP to address the Universal Domain Adaptation (UniDA) problem through label space alignment. TLSA utilizes a generative vision-language model to discover labels for all target domain samples and filters out those classified as source classes. On a mini-batch level, we eliminate lexical synonyms of source labels through synonym label alignment and remove semantically ambiguous labels via semantic label alignment. At the dataset level, frequency-based noisy candidate filtering removes unreliable labels based on occurrence patterns. This approach is highly efficient in both time and memory, completing the entire process in a single epoch without requiring further tuning. After alignment, we extend the source class set with the newly predicted target-private labels to construct a universal classifier, which significantly outperforms existing methods. Additionally, through self-training with balanced top-k confident sample selection, we surpass state-of-the-art UniDA baselines.
Data Availability
All the data utilized in this paper is publicly available at (Saenko et al., 2010; Venkateswara et al., 2017; Peng et al., 2017, 2019).
Declaration of generative AI in scientific writing
During the preparation of this work, the author used ChatGPT (Achiam et al., 2023) in order to improve readability. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
Appendix A Pseudo Code
@@ Appendix B Examples of Label Space Alignment
Appendix C Details of Semantic Label Alignment
@@
Appendix D Analysis of Labels to be Added to the Frequency Bank
In Sec. 3.4, we noted that storing BLIP’s discovered label to the frequency bank is more beneficial for performance compared to storing CLIP’s top-1 prediction to the frequency bank. As shown in Tab. 13, storing CLIP’s top-1 predictions often leads to an overestimation of incorrect target private labels (e.g., overly specific or entirely incorrect classes), which in turn causes the correct target private labels (e.g., ‘train’, ‘truck’, ‘skateboard’) to be underestimated. This imbalance ultimately results in degraded model performance. In Tab. 13, the correct target private label is underestimated when using CLIP’s top-1 prediction compared to using BLIP’s discovered label.
@@ Appendix D Analysis of Labels to be Added to the Frequency Bank
Appendix E Details of Prompts to Generative VLM
This section provides details of the question prompts used for label discovery. To ensure diversity while maintaining the same meaning, we employed five different styles of prompts. Tab. 14 demonstrates the detailed examples of question prompts for generative VLM.
@@ Appendix E Details of Prompts to Generative VLM
Appendix F Time Complexity of Our Method
Fig. 12 illustrates the time complexity of our approach. Although two VLMs are employed, the framework is entirely training-free, introducing minimal overhead compared to prior works. Furthermore, it achieves faster adaptation than clustering-based methods (Chang et al., 2022).
A potential limitation lies in scalability, since generative VLM inference is required for each target-domain sample, leading to a runtime that grows linearly with dataset size. In practice, however, the cost remains manageable: even on DomainNet, the largest dataset in DomainBed, inference completed within 1.5 hours on a single GPU, demonstrating practical feasibility.
@@ Appendix F Time Complexity of Our Method
Appendix G Details of Main Results
This section details the experimental results (H-score, H3-score) for domain-to-domain. Tab. 16 through Tab. 26 offer a comprehensive comparison between baseline methods and the proposed approach using the CLIP ViT-B/16 backbone.